OpenStudy (anonymous):

15x^2-14x-8 factoring..

OpenStudy (anonymous):

15x^2 - 14x - 8. Here's how to factor: Find a, b such that: (x + a)(x + b) = Ax^2 + Bx + C; a + b = B; ab = C

OpenStudy (anonymous):